Grade 12Physical Chemistry

what is the potential of a cell containing two hydrogen electrodes,in which the anode is in contact with 10 to the power 5 M HCL and the cathode is in contact with 1000 times the concentration of HCL as that of the anode

Akshaya jeenraj
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
In  this cathode and anode boh aare hydrogen electrode 
[H^+] in cathode = 10 (power ngative 2 ) as the con of HCl is 1000 timess greater than that of anode and [H^+]  at anode is 10 (power negative 5 )
 E^0 cell = 0 as cathode and anode are htdrogen , therefore 
 EMF of cell = 0.059 / 1   log 10^-2 / 10^ -5 = 0.059 x  3 =0.177V.